you push a 24 kg box horizontally across a warehouse floor with 66 N of applied force. There is 45 N of friction acting on the sled. Determine the acceleration of the sled in metres per second squared. Enter only the number, rounded to two significant digits. *
1
Expert's answer
2020-04-24T10:52:29-0400

The net force applied to the box

"F_{\rm net}=66-45=21\:\rm N."

The acceleration of the box

"a=\frac{F_{\rm net}}{m}=\frac{21}{24}=0.88\:\rm m/s^2."